# D2ST-Adapter: Disentangled-and-Deformable Spatio-Temporal Adapter for Few-shot Action Recognition

### v3

@@ 3.3 End-to-End Adapter Tuning
The learned features by our adapter tuning framework are further used for few-shot action recognition based on the metric-based strategy. It first calculates the frame-wise $L_{2}$ distance matrix between the input query video and each class prototype derived by simply averaging the corresponding support samples. Then the distance matrix is used to calculate the matching similarity between the query and each class for prediction. Formally, the similarity between query $q$ and class prototype $c$ can be expressed as:
$s(\mathbf{F}_{q},\mathbf{F}_{c})=\mathbf{M}([\mathbf{F}_{q}^{1},\cdots,\mathbf{F}_{q}^{T}],[\mathbf{F}_{c}^{1},\cdots,\mathbf{F}_{c}^{T}]),$ (5)
where $\mathbf{F}_{q}^{i}$ and $\mathbf{F}_{c}^{j}$ denote the features of the $i$-th frame in query $q$ and the $j$-th frame in class prototype $c$ while $\mathbf{M}$ is the matching metric such as OTAM [2] and Bi-MHM [44]. The whole model can be optimized in an end-to-end manner using the cross-entropy loss.
@@ 4.1 Experimental Setup
Datasets. We conduct experiments on five standard few-shot action recognition benchmarks, including SSv2-Full [10], SSv2-Small [10], Kinetics [3], HMDB51 [19], and UCF101 [33]. The temporal features are crucial to the action recognition in SSv2-Full and SSv2-Small while the actions in other datasets relies more on the spatial features than the dynamic features. Following the typical data split [59, 2, 43], For SSv2-Full, SSv2-Small and Kinetics, we select 64/24/12 classes from the datasets as the training/validation/test set, respectively. As for HMDB51 and UCF101, we adopt the same data split as [54, 43].
Pre-trained backbone. Most prior arts employs the ResNet-50 [13] pre-trained on ImageNet [5] as backbone to extract frame features. Recently, CLIP-FSAR [42] conduct experiments on CLIP-ViT-B [7] and get remarkable performance. To verify the effectiveness of our D2ST-Adapter, we instantiate our method with both pre-trained backbones.
Implementation details. We insert one D2ST-Adapter module in each stage of the pre-trained model (4 stages for ResNet-50 and 12 stages for CLIP-ViT-B) as shown in Figure 2 (a). We tune the sampling kernel for aDSTA-S and aDSTA-T for each backbone and the details are provided in the supplementary material. For a fair comparison with previous methods [2, 44, 43], we uniformly sample 8 frames (i.e., $T=8$) as the input of a video. We employ several standard techniques for data augmentation such as random crop and color jitter in the training stage. Bi-MHM [44] is used as the matching metric in comparison with other methods. Besides, OTAM [2] and TRX [28] are additionally used in the ablation study. We train the model episodically with Adam [18] optimizer. Following the common practice [44, 43], in the test stage, we randomly select 10,000 episodes from the test set and report the average accuracy.
@@
Method Backbone Adaptation SSv2-Full SSv2-Small
paradigm 1-shot 2-shot 3-shot 4-shot 5-shot 1-shot 3-shot 5-shot
CMN [59] ResNet-50 Full fine-tuning 34.4 -- -- -- 43.8 33.4 42.5 46.5
CMN-J [60] ResNet-50 Full fine-tuning -- -- -- -- -- 36.2 44.6 48.8
OTAM [2] ResNet-50 Full fine-tuning 42.8 49.1 51.5 52.0 52.3 36.4 45.9 48.0
ITANet [56] ResNet-50 Full fine-tuning 49.2 55.5 59.1 61.0 62.3 39.8 49.4 53.7
TRX [28] ResNet-50 Full fine-tuning 42.0 53.1 57.6 61.1 64.6 36.0 51.9 59.1
TA2[22] ResNet-50 Full fine-tuning 47.6 -- -- -- 61.0 -- -- --
MTFAN [46] ResNet-50 Full fine-tuning 45.7 -- -- -- 60.4 -- -- --
STRM [36] ResNet-50 Full fine-tuning 43.1 53.3 59.1 61.7 68.1 37.1 49.2 55.3
HyRSM [44] ResNet-50 Full fine-tuning 54.3 62.2 65.1 67.9 69.0 40.6 52.3 56.1
Nguyen et al. [24] ResNet-50 Full fine-tuning 43.8 -- -- -- 61.1 -- -- --
Huang et al. [15] ResNet-50 Full fine-tuning 49.3 -- -- -- 66.7 38.9 -- 61.6
HCL [57] ResNet-50 Full fine-tuning 47.3 54.5 59.0 62.4 64.9 38.7 49.1 55.4
SloshNet [49] ResNet-50 Full fine-tuning 46.5 -- -- -- 68.3 -- -- --
GgHM [50] ResNet-50 Full fine-tuning 54.5 -- -- -- 69.2 -- -- --
MoLo [43] ResNet-50 Full fine-tuning 56.6 62.3 67.0 68.5 70.6 42.7 52.9 56.4
ST-Adapter [25] ResNet-50 Adapter tuning 52.2 59.9 64.1 67.1 68.7 41.9 51.5 55.7
D2ST-Adapter (Ours) ResNet-50 Adapter tuning 57.0 65.2 69.5 71.4 73.6 45.8 56.6 60.9
CLIP-FSAR [42] CLIP-ViT-B Full fine-tuning 61.9 64.9 68.1 70.9 72.1 54.5 58.6 61.8
AIM [51] CLIP-ViT-B Adapter tuning 63.7 72.0 75.2 77.3 79.2 52.8 64.0 67.5
DUALPATH [26] CLIP-ViT-B Adapter tuning 64.5 72.6 76.0 78.0 79.8 53.5 64.4 68.1
ST-Adapter [25] CLIP-ViT-B Adapter tuning 64.2 72.4 76.0 77.4 79.5 53.1 64.2 68.0
D2ST-Adapter (Ours) CLIP-ViT-B Adapter tuning 66.7 75.3 78.3 80.1 81.9 55.0 65.8 69.3
@@
4.2 Comparison with State-of-the-Art Methods
Benchmarks sensitive to temporal features. Table 1 shows the comparative results of different methods on SSv2-Full and SSv2-Small, which are quite challenging since both the temporal and spatial features are crucial to the action recognition. Our D2ST-Adapter achieves the best performance and outperforms other methods by a large margin with both backbones in all few-shot settings except ‘5-shot’ on SSv2-Small. These results demonstrate the advantages of our method over other methods in dealing with challenging scenarios where the temporal features are critical for action recognition. In particular, our D2ST-Adapter outperforms all existing adapter tuning methods for video data, including AIM, DUALPATH, and ST-Adapter, substantially on all settings, which reveals the benefit of two core technical designs of D2ST-Adapter, namely the dual-pathway architecture for disentangled learning of spatial and temporal features and anisotropic Deformable Spatio-Temporal Attention (aDSTA) for learning features in a global view.
@@ 4.2 Comparison with State-of-the-Art Methods
To gain more insight into the strengths and weaknesses of our D2ST-Adapter, we select four categories of actions where our method excels, as well as four categories where our model struggles in comparison to the state-of-the-art MoLo and ST-Adapter.  Figure 3 shows that our model achieves large performance superiority over other methods on the relatively complex actions requiring careful reasoning via learning temporal features for recognition, whilst performing poorly on simple actions that can be recognized primarily based on spatial features. Such results are consistent with the previous results that our model excels at recognizing actions involving complex temporal dynamics.
[…]
Method Backbone Adaptation Kinetics HMDB51 UCF101
paradigm 1-shot 2-shot 3-shot 4-shot 5-shot 1-shot 3-shot 5-shot 1-shot 3-shot 5-shot
CMN [59] ResNet-50 Full fine-tuning 57.3 67.5 72.5 74.7 76.0 -- -- -- -- -- --
CMN-J [60] ResNet-50 Full fine-tuning 60.5 70.0 75.6 77.3 78.9 -- -- -- -- -- --
OTAM [2] ResNet-50 Full fine-tuning 73.0 75.9 78.7 81.9 85.8 54.5 65.7 68.0 79.9 87.0 88.9
ITANet [56] ResNet-50 Full fine-tuning 73.6 -- -- -- 84.3 -- -- -- -- -- --
TRX [28] ResNet-50 Full fine-tuning 63.6 76.2 81.8 83.4 85.9 53.1 66.8 75.6 78.2 92.4 96.1
TA2[22] ResNet-50 Full fine-tuning 72.8 -- -- -- 85.8 59.7 -- 73.9 81.9 -- 95.1
MTFAN [46] ResNet-50 Full fine-tuning 74.6 -- -- -- 87.4 59.0 -- 74.6 84.8 -- 95.1
STRM [36] ResNet-50 Full fine-tuning 62.9 76.4 81.1 83.8 86.7 52.3 67.4 77.3 80.5 92.7 96.9
HyRSM [44] ResNet-50 Full fine-tuning 73.7 80.0 83.5 84.6 86.1 60.3 71.7 76.0 83.9 93.0 94.7
Nguyen et al. [24] ResNet-50 Full fine-tuning 74.3 -- -- -- 87.4 59.6 -- 76.9 84.9 -- 95.9
Huang et al. [15] ResNet-50 Full fine-tuning 73.3 -- -- -- 86.4 60.1 -- 77.0 71.4 -- 91.0
HCL [57] ResNet-50 Full fine-tuning 73.7 79.1 82.4 84.0 85.8 59.1 71.2 76.3 82.5 91.0 93.9
SloshNet [49] ResNet-50 Full fine-tuning 70.4 -- -- -- 87.0 59.4 -- 77.5 86.0 -- 97.1
GgHM [50] ResNet-50 Full fine-tuning 74.9 -- -- -- 87.4 61.2 -- 76.9 85.2 -- 96.3
MoLo [43] ResNet-50 Full fine-tuning 74.0 80.4 83.7 84.7 85.6 60.8 72.0 77.4 86.0 93.5 95.5
ST-Adapter [25] ResNet-50 Adapter tuning 73.0 79.9 82.8 84.6 85.1 60.3 71.4 74.7 84.6 92.9 94.5
D2ST-Adapter (Ours) ResNet-50 Adapter tuning 75.0 81.4 84.3 85.8 87.0 61.6 73.0 76.6 86.9 94.4 95.6
CLIP-FSAR [42] CLIP-ViT-B Full fine-tuning 89.7 92.9 94.2 94.8 95.0 75.8 84.1 87.7 96.6 98.4 99.0
AIM [51] CLIP-ViT-B Adapter tuning 88.4 92.7 94.2 94.6 95.3 74.2 83.8 86.9 95.4 98.2 98.5
DUALPATH [26] CLIP-ViT-B Adapter tuning 88.8 92.8 94.3 94.7 95.4 74.9 84.6 87.5 95.7 98.4 98.7
ST-Adapter [25] CLIP-ViT-B Adapter tuning 88.5 92.6 94.0 94.6 95.1 74.1 84.3 87.3 95.9 98.5 98.9
D2ST-Adapter (Ours) CLIP-ViT-B Adapter tuning 89.3 93.1 94.4 95.2 95.5 77.1 86.4 88.2 96.4 98.7 99.1
Benchmarks relying on spatial features. Table 2 presents the experimental results on the benchmarks relying more on spatial features, including Kinetics, HMDB51 and UCF101. Our method still performs best on most of the settings, which manifest the effectiveness and robustness of our method in these scenarios. We also observe that the performance improvement by our D2ST-Adapter over other method, especially ST-Adapter, is smaller than that on SSv2 datasets. It is reasonable since the action recognition in these datasets relies less on the temporal features and thus the disentangled feature encoding of our method yields limited performance gain.
@@
Comparison of efficiency. We compare the efficiency between our D2ST-Adapter and other methods in terms of tunable parameter size and memory usage in  Figure 4. It shows that the tunable parameter size of both our model and ST-Adapter is significantly smaller than other methods that are based on full fine-tuning, which reveals the advantage of the adapter tuning paradigm over the full fine-tuning paradigm for task adaptation. Meanwhile, our method outperforms these full fine-tuning based methods substantially while enjoying much fewer tunable parameters, which validates the superiority of our method.
4.3 Ablation study
Comparison between different Adapters. We first compare the performance of different Adapters in Table 3. As shown in Figure 1, Vanilla-Adapter [14] only contains the downsampling and upsampling layers with a GELU nonlinearity in between. As described in Section 3.2, DST-Adapter is the convolutional version of our D2ST-Adapter which follows the dual-pathway structure but models spatial and temporal pathways with 3D convolution. The performance of full fine-tuning is also provided for reference. We make following observations.
  • Effect of disentangled encoding of the spatial and temporal features. The performance improvement from ST-Adapter to DST-Adapter demonstrates the effectiveness of dual-pathway adapter architecture, which allows for encoding the spatial and temporal features in a disentangled manner.
  • Effect of the proposed anisotropic Deformable Spatial-Temporal Attention (aDSTA). Comparing our D2ST-Adapter with DST-Adapter, the proposed aDSTA yields a large performance gain, especially on the SSv2-Full dataset, which is more challenging due to the necessity of encoding temporal features.
Table 3: Comparison between different adapters as well as full fing-tuning.
@@ 4.3 Ablation study
DST-Adapter (Ours) 53.7 69.6 73.3 85.5
D2ST-Adapter (Ours) 57.0 73.6 75.0 87.0
Table 4: Ablation study on the configuration of sampling kernel of aDSTA.
Configuration SSv2-Full Kinetics
@@
aDSTA-S & aDSTA-T 57.0 73.6 75.0 87.0

### v4

@@ 3.3 End-to-End Adapter Tuning
The learned features by our adapter tuning framework are further used for few-shot action recognition based on the metric-based strategy. Following ProtoNet [26], it first calculates the frame-wise $L_{2}$ distance matrix between the query video and each class prototype derived by averaging the corresponding support samples. Then the distance matrix is used to calculate the matching similarity between the query and each class for prediction. Formally, the similarity between query $q$ and class prototype $c$ can be expressed as:
$\vspace{-3pt}s(\mathbf{F}_{q},\mathbf{F}_{c})=\mathbf{M}([\mathbf{F}_{q}^{1},\cdots,\mathbf{F}_{q}^{T}],[\mathbf{F}_{c}^{1},\cdots,\mathbf{F}_{c}^{T}]),$ (5)
where $\mathbf{F}_{q}^{i}$ and $\mathbf{F}_{c}^{j}$ denote the features of the $i$-th frame in query $q$ and the $j$-th frame in class prototype $c$, and $\mathbf{M}$ is the matching metric such as Bi-MHM [33]. During training, we take the similarities for each class as logits and optimize the model in an end-to-end manner using the same cross-entropy loss as in [3, 23, 33]. For inference, we classify the query as the support class closest to it.
@@ 4.1 Experimental Setup
Datasets. We conduct experiments on five standard few-shot action recognition benchmarks, including SSv2-Full [8], SSv2-Small [8], Kinetics [14], HMDB51 [15], and UCF101 [27]. Following the typical data split [51, 3, 34], for SSv2-Full, SSv2-Small and Kinetics, we select 64/12/24 classes from the datasets as the training/validation/test set, respectively. As for HMDB51 and UCF101, we adopt the same data split as [45, 34].
@@
Method Pre-training Fine-tuning SSv2-Full SSv2-Small Kinetics HMDB51 UCF101
1-shot 5-shot 1-shot 5-shot 1-shot 5-shot 1-shot 5-shot 1-shot 5-shot
CMN [51] INet-RN50 Full 34.4 43.8 33.4 46.5 57.3 76.0 -- -- -- --
OTAM [3] INet-RN50 Full 42.8 52.3 36.4 48.0 73.0 85.8 54.5 68.0 79.9 88.9
AMeFu-Net [7] INet-RN50 Full -- -- -- -- 74.1 86.8 60.2 75.5 85.1 95.5
ITANet [46] INet-RN50 Full 49.2 62.3 39.8 53.7 73.6 84.3 -- -- -- --
TRX [23] INet-RN50 Full 42.0 64.6 36.0 59.1 63.6 85.9 53.1 75.6 78.2 96.1
TA2[18] INet-RN50 Full 47.6 61.0 -- -- 72.8 85.8 59.7 73.9 81.9 95.1
MTFAN [38] INet-RN50 Full 45.7 60.4 -- -- 74.6 87.4 59.0 74.6 84.8 95.1
STRM [28] INet-RN50 Full 43.1 68.1 37.1 55.3 62.9 86.7 52.3 77.3 80.5 96.9
HyRSM [33] INet-RN50 Full 54.3 69.0 40.6 56.1 73.7 86.1 60.3 76.0 83.9 94.7
Nguyen et al. [20] INet-RN50 Full 43.8 61.1 -- -- 74.3 87.4 59.6 76.9 84.9 95.9
Huang et al. [12] INet-RN50 Full 49.3 66.7 38.9 61.6 73.3 86.4 60.1 77.0 71.4 91.0
HCL [48] INet-RN50 Full 47.3 64.9 38.7 55.4 73.7 85.8 59.1 76.3 82.5 93.9
SA-CT [47] INet-RN50 Full 48.9 69.1 -- -- 71.9 87.1 60.4 78.3 85.4 96.4
SloshNet [42] INet-RN50 Full 46.5 68.3 -- -- 70.4 87.0 59.4 77.5 86.0 97.1
GgHM [43] INet-RN50 Full 54.5 69.2 -- -- 74.9 87.4 61.2 76.9 85.2 96.3
MoLo [34] INet-RN50 Full 56.6 70.6 42.7 56.4 74.0 85.6 60.8 77.4 86.0 95.5
Zheng et al. [49] INet-RN50 Full 55.4 71.4 41.3 57.5 75.6 87.6 61.5 76.0 87.8 96.0
ST-Adapter [21] INet-RN50 PEFT 52.2 68.7 41.9 55.7 73.0 85.1 60.3 74.7 84.6 94.5
D2ST-Adapter INet-RN50 PEFT 57.0 73.6 45.8 60.9 75.8 87.7 61.6 76.6 86.9 95.6
AIM [44] CLIP-ViT-B PEFT 63.7 79.2 52.8 67.5 88.4 95.3 74.2 86.9 95.4 98.5
DUALPATH [22] CLIP-ViT-B PEFT 64.5 79.8 53.5 68.1 88.8 95.4 74.9 87.5 95.7 98.7
ST-Adapter [21] CLIP-ViT-B PEFT 64.2 79.5 53.1 68.0 88.5 95.1 74.1 87.3 95.9 98.9
D2ST-Adapter CLIP-ViT-B PEFT 66.7 81.9 55.0 69.3 89.3 95.5 77.1 88.2 96.4 99.1
CLIP-FSAR [35] CLIP-ViT-B Full 62.1 72.1 54.6 61.8 94.8 95.4 77.1 87.7 97.0 99.1
EMP-Net [37] CLIP-ViT-B PEFT 63.1 73.0 57.1 65.7 89.1 93.5 76.8 85.8 94.3 98.2
MA-FSAR [41] CLIP-ViT-B PEFT 63.3 72.3 59.1 64.5 95.7 96.0 83.4 87.9 97.2 99.2
Task-Adapter [2] CLIP-ViT-B PEFT 71.3 74.2 60.2 70.2 95.0 96.8 83.6 88.8 98.0 99.4
D2ST-Adapter-MM CLIP-ViT-B PEFT 74.1 83.6 63.6 71.9 96.2 97.2 84.3 89.0 98.3 99.5
Implementation details. To evaluate D2ST-Adapter, we instantiate our method with both ResNet-50 [10] and CLIP-ViT-B/16 [5]. We insert one D2ST-Adapter module in each stage of the pre-trained model (4 stages for ResNet-50 and 12 stages for CLIP-ViT-B/16), as shown in Figure 2 (a). For a fair comparison with previous methods [33, 34], we uniformly sample 8 frames (i.e. $T=8$) as the input of a video. Bi-MHM [33] is used as the matching metric in comparison with state-of-the-art methods. Besides, OTAM [3] and TRX [23] are additionally employed in the ablation study. We report the average classification accuracy of 10,000 episodes randomly selected from the test set. Other implementation details follow HyRSM [33], which are provided in the supplementary material.
4.2 Comparison with State-of-the-Art Methods
We compare our method with other methods on two types of datasets: temporal-related benchmarks including SSv2-Full and SSv2-Small datasets, and spatial-related benchmarks inlcuding Kinetics, HMDB51, and UCF101 datasets.
Benchmarks sensitive to temporal features. Table 1 shows the comparative results of different methods on two temporal-related datasets, i.e. SSv2-Full and SSv2-Small, which are quite challenging since both the temporal and spatial features are crucial to the action recognition. Our D2ST-Adapter achieves the best performance and outperforms other methods by a large margin in most few-shot settings. These results demonstrate the advantages of our method over other methods in dealing with challenging scenarios where the temporal features are critical for action recognition. In particular, our D2ST-Adapter substantially outperforms all adapter tuning methods for video data including AIM, DUALPATH, and ST-Adapter, highlighting the advantages of its core design, namely built-in disentangled adaptation of spatial and temporal features.
Benchmarks relying on spatial features. Experimental results on the spatial-related benchmarks including Kinetics, HMDB51, and UCF101, are also presented in Table 1. Our method still performs best in most settings, which manifest the effectiveness and robustness of our method in these scenarios. Besides, the performance improvement by our D2ST-Adapter over other method, especially ST-Adapter, is smaller than that on SSv2 datasets. It is reasonable since the action recognition in these datasets relies less on the temporal features and thus the disentangled feature encoding of our method yields limited performance gain.
Table 2: Comparison with previous adapter-based methods on traditional full-shot action recognition task with CLIP-ViT-B/16 backbone. The testing views (# frames per clip $\times$ # temporal clip $\times$ # spatial crop) are set as 8$\times$1$\times$3 for K400 and 8$\times$3$\times$1 for SSv2.
Method K400 SSv2
AIM [44] 83.9 66.4
DUALPATH [22] 84.1 67.3
ST-Adapter [21] 82.0 67.1
D2ST-Adapter (Ours) 84.5 68.0
Comparison with multimodal methods. To have a fair comparison with the multimodal methods for few-shot action recognition, we extend our model to multi-modality (D2ST-Adapter-MM). This extension draws on the multimodal framework of CLIP-FSAR [35] to align with other methods. Table 1 shows that our method outperforms all other multimodal methods, especially on SSv2-Full and SSv2-Small datasets. These findings consistently highlight the effectiveness and robustness of our D2ST-Adapter.
@@ 4.2 Comparison with State-of-the-Art Methods
Comparison on traditional (full-shot) action recognition task. Our model is particularly effective in the few-shot scenario while it can also be applied to traditional action recognition tasks using full training samples. To obtain a more comprehensive assessment of our model, we conduct experiments on two traditional action recognition benchmarks, Kinetics-400 (K400) [14] and Something-Something-v2 (SSv2) [8], using the same settings as ST-Adapter [21]. Table 2 shows that our method still outperforms these adapter-based methods on both benchmarks.
@@
Case Study of the Pros and Cons of D2ST-Adapter. To gain more insights into the strengths and weaknesses of our D2ST-Adapter, we select four categories of actions where our method excels at, as well as four categories where our model struggles in comparison to the state-of-the-art MoLo and ST-Adapter. Figure 3 shows that our model achieves large performance superiority over other methods on the relatively complex actions requiring careful reasoning via learning temporal features for recognition, whilst performing on par with other methods on simple actions that can be recognized primarily based on spatial features. Such results are consistent with the conclusion that our model excels at recognizing actions involving complex temporal dynamics.
Table 3: Comparison with other adapter-based methods using CLIP-ViT-B/16 as backbone on efficiency. Accuracy on SSv2-Full benchmark in 1-shot setting is also provided.
Method Tunable Memory GFLOPs Latency Acc.
Params Usage
AIM [44] 14.3 M 19.0 GB 208 14.1 ms 63.7
DUALPATH [22] 14.3 M 16.7 GB 171 11.6 ms 64.5
ST-Adapter [21] 14.3 M 16.1 GB 163 12.6 ms 64.2
D2ST-Adapter 7.3 M 17.3 GB 150 18.2 ms 66.7
@@
Comparison of efficiency. We compare the efficiency between our D2ST-Adapter and other methods in terms of tunable parameter size and memory usage in Figure 4. It shows that the tunable parameter size of both our model and ST-Adapter is significantly smaller than other methods that are based on full fine-tuning, which reveals the advantage of the adapter tuning paradigm over the full fine-tuning paradigm for task adaptation. Meanwhile, our method outperforms these full fine-tuning based methods substantially with much lower parameter-finetuning overhead, which validates the superiority of our method.
We further compare our model with other adapter-based methods on efficiency in Table 3. Our D2ST-Adapter performs best while maintaining relatively high efficiency due to the lightweight design of the proposed aDSTA module and the smaller number of inserted adapters (only one per block in our method vs. two or four per block in other adapter-based methods [21, 44, 22]).
Figure 5: Visualization of important shifted reference points w.r.t all queries (first sample), and important shifted reference points w.r.t a given query, marked with a red star (second sample). More visualization results are provided in the supplementary material.
Table 4: Comparison between different adapters.
@@
DST-Adapter (Ours) 53.7 69.6 74.4 86.4
D2ST-Adapter (Ours) 57.0 73.6 75.8 87.7
Table 5: Effect of the anisotropic sampling of aDSTA.
Method Configurations SSv2-Full Kinetics
@@
Dual branches aDSTA-S & aDSTA-T 57.0 73.6 75.8 87.7
4.3 Ablation study
Comparison between different Adapters. We first compare the performance of different Adapters in Table 4. Vanilla-Adapter [11] only contains the downsampling and upsampling layers with a GELU nonlinearity in between. As described in Section ‘Method’, DST-Adapter is the convolutional version of our D2ST-Adapter which models both the spatial and temporal pathways with 3D convolution. The performance of full fine-tuning is also provided for reference. We make following observations.
  • Effect of disentangled encoding of the spatial and temporal features is demonstrated by the performance improvement from ST-Adapter to DST-Adapter benefited from the dual-pathway adapter architecture.
  • Effect of the proposed aDSTA. Comparing our D2ST-Adapter with DST-Adapter, the proposed aDSTA yields a large performance gain, especially on SSv2-Full which demands effective encoding of temporal features.
Effect of anisotropic sampling of aDSTA. To investigate the effect of proposed anisotropic sampling scheme of aDSTA, we compare the performance of our D2ST-Adapter using four different configurations of the aDSTA: 1) only ‘aDSTA-S’ modeling the spatial pathway is used; 2) only ‘aDSTA-T’ for the temporal pathway is preserved; 3) ‘aDSTA-Uniform’ performs sampling with uniform density in both the spatial and temporal domains; and 4) ‘aDSTA-S & aDSTA-T’ corresponds the fully functional D2ST-Adapter. Table 5 shows that the distinct performance gap between these two settings reveals the advantage of configuring specialized sampling densities for the spatial and temporal pathways.